FEL Upgrade Project Weekly Brief
July 24-28, 2000


The Senate on Thursday passed the FY2001 Defense Appropriations bill (passed by the House last week).  This bill includes $5M of funding for ONR which will hopefully be directed to the completion of the FEL
Upgrade project (after appropriate taxes and negotiation of an amended DOE/DON agreement).

We re-established generation of short pulse (sub-picosecond) x-rays this week by Compton scattering of the IR Demo IR pulses off of the driver electron beam.  The experimental configuration has been enhanced to
allow further characterization of the x-ray flux.


The FEL management team worked with Accelerator Division management to ensure that we had our best labor estimates in hand for all the required work that is sharing lab design resources (the FEL Upgrade, SNS and the CEBAF Upgrade).  We are prioritizing the scheduled work and making sure we maintain flexibility in sharing the available resources.

Two papers were submitted to the Directed Energy Symposium which will be held the first week in November in Albuquerque.

Design/Construction Activities

WBS 3 (Beam Physics)

The beam physics group produced a draft paper for the Linac Conference which describes the present design of the accelerator driver.

WBS 4 (Injector)

Concerning our test of the Apertured Cesiator:  we have installed a residual gas analyzer and new ion pump on the system and are analyzing results.  With the system cold we note some hydrocarbon contamination.
We are replacing a suspect component and will then purge the system with dry, warm nitrogen before proceeding with the subsequent vacuum qualification.

We received a rough design for a low energy cesium ion gun from a vendor that may present another option for directed cesiation.  However, we doubt that the zeolite source, which is used in these commercial sources, can be made pure enough for our standards.


Following last week's report that noted the progress in the two test cavities that will be shipped to KEK to test electropolishing:  the first cavity (which successfully reached 14MV/m after our "standard" chemistry) is ready for shipment to KEK.  The second cavity has completed only the initial chemistry and is presently under going tuning prior to final chemistry and installation in  the VTA dewar for cold testing.

WBS 6 (RF)

We installed waveguide from the second floor power systems location through the penetrations to the FEL accelerator vault for powering the number two cryomodule.  In addition, the second High Power Amplifier
(HPA) for powering the second cryomodule klystrons was moved into position.

WBS 8  (I & C)

The drawings for the 3" beam viewers were delayed one week due to shifting manpower.  These should be ready for checking next week.  The final count will be firmed up in August.  We are getting 10 of the
frames built for test fit-up with the 1.5 micron Al foil.  We continued to work on 2" vs. 3" hardware cost/benefit estimates.

WBS 9 (Transport)

o We continued layout of the GX Prototype concentrating on the cooling water manifold and connections, the internal bus bar connections and the external lead connections.
o For the 3D model of the GX dipole, the steel is entered into the code successfully but the coil looks to the program like a discontinuous object.  We are looking for an entry error.

o Pre finish machining of one steel quadrant is due August 4th with CNC machining starting on Monday, August 7th.
o After small improvements to the winding fixture at startup, winding of the first coil has restarted.

o A small part of the time of the two engineers diverted to other duties has been retained to keep the ongoing projects moving through development.
o Proposals for the engineering services contract continue in evaluation.
o New alignment cartridges are out for bid.

WBS 10 (Wiggler)

o Core modifications for the Optical Klystron are completed.  Design of the diagnostics, vacuum pipe, and water manifolds continues.

o Details of all parts for the Dispersion Section are ready for sign off and the requisite steel has been delivered to the shop for start of fabrication.

WBS 11 (Optics)

We submitted a requisition to continue the design efforts for the figure scheme that was generated in our initial design work package.


After a very successful run so far this month in terms of the operational reliability of the FEL, this week's operations were problematic.  We had problems with: (1) one cavity mirror (braze failure in the mount for the 90% reflector for a 5 micron cavity mirror installed last week end); (2) failure of a trim magnet power supply
which was installed to replace a trim power supply which was showing sporadic problems; (3) various problems with phase locking of the drive laser which were partially traced to a failed coaxial relay in the laser
manufacturer's mode locking circuitry, and (4) overheating of one of the HPA's traced to a carbonized contact.

Despite these problems, we have delivered light to the Vanderbilt-CWM team to continue their highly successful Si spectroscopy experiments.  (The results from their initial studies using the IR Demo will be
published in Physical Review Letters next month.)  Also on the owl shift we set up the second generation experiment (J. Boyce, G. Krafft, et al.) to detect and characterize short pulse x-rays produced by Compton
scattering.  The near term goal was to better characterize the energy and spatial distribution of the x-rays so that hardware could be designed to allow x-ray/IR pump probe experiments.

The I&C team performed the following services for users:

The time spent setting up and diagnosing the Compton x-ray experiment was worthwhile. The data taking went well with the observation of 100:1 SNR as seen before. We had hoped to use a commercial x-ray camera to map the x-ray distribution but the camera was washed out by background so a pin diode detector will be mounted to a translation stage to map the distribution. We performed a noise-reduction exorcism in lab 1 for the Vanderbilt-CWM Si spectroscopy experiments.  The experimenters tend to pile all of their equipment together; preamp & stepper motor drivers!

In addition we coordinated installation of additional AC power of lab 3 for the new carbon nanotube experiments (PSU-CWM-NASA), and also rigged in a TC controller to replace the blown one from their carbon oven.

For the remaining week of the user run, Bob Austin from Princeton University is scheduled for a second run with protein spectroscopy at 6 microns and in addition we will attempt to get some light to the nanotube group in a parasitic fashion so that they can begin to test their hardware. A special weekend run is planned this last weekend of the run to continue the x-ray characterization.

Privacy and Security Notice